OpenStudy (anonymous):

Find the rate of change between the two points: (4, 10) and (6, 15) where x is in minutes and y is in miles. A. 4 miles per minute B. 4 minutes per mile C. 2.5 miles per minute D. 2.5 minutes per mile will fan and medal

myininaya (myininaya):

use the slope formula

myininaya (myininaya):

\[\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

no it didnt thank you though

myininaya (myininaya):

baseball just enter the points in that formula above \[x_1=4 \\y_1=10 \\x_2=6 \\y_2=15\]

OpenStudy (anonymous):

So its D ?

OpenStudy (anonymous):

@myininaya ?

myininaya (myininaya):

well it says y is in miles and x is in minutes

myininaya (myininaya):

didn't we have y's on top?

OpenStudy (anonymous):

ya i also think it is D!

OpenStudy (anonymous):

^^^

myininaya (myininaya):

It isn't D.

OpenStudy (anonymous):

C ? @myininaya

myininaya (myininaya):

we have change of y /change of x and y is in miles while x is in minutes so miles should be over minutes

OpenStudy (anonymous):

A @ myininaya

myininaya (myininaya):

what know you just said we had 2.5 i'm telling you what the units should be

OpenStudy (anonymous):

i dont get it

myininaya (myininaya):

you did the calculation and you got 2.5 you did that by calculating change of y over change of x the top had units miles while the bottom had units minutes that is miles per minutes since miles is on top while minutes is on bottom

OpenStudy (anonymous):

So it is C

myininaya (myininaya):

\[\frac{15 miles -10 miles}{6 \min-4 \min}=\frac{5 miles}{2 \min} =2.5 \frac{miles}{\min}\] yes you read this as 2.5 miles per min

OpenStudy (anonymous):

thank you